# FLUID: A Fine-Grained Lightweight Urban Signalized-Intersection Dataset of Dense Conflict Trajectories

### v1

@@ Abstract
The trajectory data of traffic participants (TPs) is a fundamental resource for evaluating traffic conditions and optimizing policies, especially at urban intersections. Although data acquisition using drones is efficient, existing datasets still have limitations in scene representativeness, information richness, and data fidelity. This study introduces FLUID, comprising a fine-grained trajectory dataset that captures dense conflicts at typical urban signalized intersections, and a lightweight, full-pipeline framework for drone-based trajectory processing. FLUID covers three distinct intersection types, with approximately 5 hours of recording time and featuring over 20,000 TPs across 8 categories. Notably, the dataset averages two vehicle conflicts per minute, involving roughly 25% of all motor vehicles. FLUID provides comprehensive data, including trajectories, traffic signals, maps, and raw videos. Comparison with the DataFromSky platform and ground-truth measurements validates its high spatio-temporal accuracy. Through a detailed classification of motor vehicle conflicts and violations, FLUID reveals a diversity of interactive behaviors, demonstrating its value for human preference mining, traffic behavior modeling, and autonomous driving research.
@@
Methods for observing traffic participant (TP) mobility and interactions are primarily categorized into ego-centric, roadside, and aerial perspectives [7]. Ego-centric approaches, foundational to autonomous driving datasets like KITTI [8], Apolloscape [9], Waymo [10], and nuPlan [11], equip vehicles with multi-modal sensors to capture surrounding interactions [12], but are inherently limited by occlusions and a restricted field of view that prevent a complete scene understanding. Roadside methods, employing either single-sensor (e.g., NGSIM [13], Zen Traffic Data [14], I-24 MOTION [15], TJRD [16]) or multi-sensor systems (e.g., DLR-UT [17]), offer long-term observation but suffer from limited spatial coverage and potential error accumulation [18]. A critical drawback for both ground-based methods is that visible hardware can alter driver behavior, compromising data naturalness [19]. In contrast, aerial observation—typically using drones for their stability and cost-effectiveness—naturally overcomes these limitations by providing a complete, bird’s-eye view of the entire traffic scene and all simultaneous interactions within it. Crucially, high-altitude operation ensures the drone remains virtually unnoticed, thus preserving the naturalness of TP behavior. These advantages have spurred the development of numerous drone-based datasets for highways (e.g., highD [20], exiD [21], CQSkyEyeX [22], MiTra [23]), urban junctions (e.g., inD [24], SIND [25], Hohhot-HDI [26], RounD [27]), and mixed urban environments (e.g., INTERACTION [28], pNEUMA [29], CitySim [30]).
Among the various traffic scenarios, urban intersections are particularly critical due to the significant efficiency losses and safety risks associated with the interrupted traffic flow. In the domain of intersection datasets, our review focuses on three key aspects: the captured scenes, the provided information, and the quality of trajectory data. Prominent existing datasets typically cover a limited number of scenes (e.g., 3-4 intersections) [28, 24, 30, 25], but generally provide comprehensive information, including precise object dimensions and detailed scene elements. In contrast, the HDI dataset [26], while featuring diverse trajectory patterns across more intersections, lacks the aforementioned details. This discrepancy precludes a direct and fair comparison. Therefore, our subsequent analysis evaluates the datasets presented in Table 1 against these three aspects, with separate considerations for motor vehicles (MVs) and vulnerable road users (VRUs, including pedestrians and two-wheelers).
@@
Geometries and channelization Control methods MV arrival rate ($min^{-1}$) VRU arrival rate ($min^{-1}$) Conflict MV ratio Assoc. MV ($\leq$10 m) MV types VRU types Conflict annot. Intention annot. Speed Position
INTERACTION 2019 FI, TI U (AWS) 34 2 1.81% 0.94 car pedestrian / bicycle
inD 2019 FI, TI U (RBL/prio.) 13 5 5.70% 1.01 car pedestrian, bicycle
CitySim 2022 FI, TI S (prot./perm.) / U (prio.) 25 18.60% 1.43 car
SIND 2022 FI S (perm.) 21 12 13.55% 1.59 car, truck, bus pedestrian, bicycle, motorcycle $\checkmark$
FLUID (ours) 2025 FI, TI, FIDRT S (prot./perm.) 30 37 24.85% 1.58 car, tricycle, van, truck, bus, trailer pedestrian, moped $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
  • INTERACTION dataset: We use the publicly available multi-agent prediction version, which features pre-extracted interaction scenarios.
  • FI/TI/FIDRT: Abbreviations for intersection geometry and channelization types, as defined in the text.
@@
  • MV/VRU arrival rate: The average number of new MVs/VRUs arriving per minute during the recording period.
@@
The first aspect is the captured scenes, analyzed from both static (intersection type) and dynamic (traffic flow) perspectives.From a static perspective, intersections are classified by their geometries and channelization—such as three-way/four-way intersections (TI/FI), and four-way with dedicated right-turn lanes (FIDRT)—and by their control methods. The latter include unsignalized (e.g., uncontrolled/right-before-left, all-way-stop, priority-controlled) and signalized types, which feature permissive or protected phases for conflicting directions [31]. Different characteristics of intersections directly affect the speed and angle of TPs when crossing them, which is closely related to the situation of traffic conflicts. However, Figure 1 shows that existing drone datasets seldom focus on intersections with dedicated channelization and usually provide incomplete coverage of diverse control strategies. From a dynamic perspective, we evaluate traffic flow using three key metrics: TP arrival rates, MV conflict ratio, and the number of associated MV per conflict ($N_{CMVCP}$), where conflicts are identified using time-based surrogate safety measures (SSMs). A comparative analysis reveals several limitations in existing datasets, such as low or imbalanced participant counts(e.g., a low VRU proportion in INTERACTION and a low overall TP count in inD) and consistently low MV conflict ratio across several major datasets. In contrast, our proposed FLUID dataset shows clear advantages, featuring a higher overall TP arrival rate and a significantly higher MV conflict ratio of over 25%. Its $N_{CMVCP}$ metric, comparable to that of SIND, further underscores its strength in capturing dense and interconnected traffic conflicts.
@@ Methods
The raw data was collected by a three-person team: one operator for the drone, and two observers who recorded traffic signal phases using ground-based cameras. We used a DJI Mini 3 drone to capture high-definition video at 4K resolution (3840×2160 pixels) and approximately 30 FPS. The drone maintained a consistent flight altitude of $100\sim 120$ meters ($100\sim 105\;m$ for the FI scene to capture finer VRU details; 120m120\;m for FIDRT and TI). The plane drift of this type of drone during stationary shooting does not exceed 1.5 meters, with a maximum recording duration per flight of up to 30 minutes. The precise flight altitude for each session can be retrieved from the drone’s flight logs. To synchronize the timestamps, the drone and ground cameras were aligned to a unified mobile phone clock with second-level precision, using the coordinated movement of a ground marker as a temporal reference.
@@
Due to the drone’s lightweight design, the raw footage was susceptible to wind-induced instability and required stabilization. We implemented a two-stage stabilization pipeline based on the open-source tools developed by Fonod et al. [34]. The lower level employed the AKAZE [35] feature detector, which offers a favorable balance between speed and performance compared to alternatives like SIFT, ORB, or BRISK, by constructing a scale space using non-linear diffusion filtering. The upper level then utilized a RANSAC algorithm with block matching and motion compensation to robustly estimate inter-frame motion parameters while rejecting false matches from dynamic objects. This process effectively eliminated significant jitter from the footage. Moreover, we develop a masking program to define a Region of Interest (ROI) that strictly encompassed the intersection area. This step served to both anonymize areas outside the road network and reduce the computational load for subsequent processing. The FIDRT scene was exempted from masking to preserve the full range of VRU activities. Finally, the videos were downsampled to 10 FPS. This frame rate is sufficient to capture the decision-making time range of human TPs and is also conducive to detecting the normal motion of VRUs [25]. The primary output of this stage is the set of pre-processed videos.
@@
Detection was accomplished using the YOLOv8 architecture [36], selected for its native support for both horizontal/oriented bounding boxes (HBB/OBB). We prioritized an OBB representation as it provides a tighter and more accurate encapsulation of object boundaries and orientation from our aerial perspective, which is critical for the subsequent analysis of kinematic parameters and TP behaviors. To overcome single-dataset limitations and achieve higher detection accuracy across diverse object classes, we employ a multi-detector ensemble strategy. This strategy involves training the same YOLOv8 architecture independently on three distinct datasets, resulting in three sets of specialised model weights, each optimised to leverage the unique strengths of its training data.
@@
For each of the three training processes, the respective dataset was re-partitioned into training, validation, and test subsets using a 70%/20%/10% ratio, and each model instance was trained for 200 epochs. The detections was then generated through a category-based fusion of the outputs from these specialised detectors, as illustrated in Figure 4. In this process, We identified the object categories where each detector outperforms the others (termed its advantage categories). The final detection output for any given frame was constructed by integrating only the detections for these designated advantage categories from their respective specialist detector. This ensemble approach ensures a comprehensive and precise set of object detections, forming a robust input for the final trajectory generation.
Lightweight Tracking.
@@
The most common time-based SSM, Post-Encroachment Time (PET), is ill-suited for this task because it requires a clear exit time from a conflict zone, which is ambiguous when considering full bounding boxes. We therefore turned to the concept of Gap Time (GT), which only considers the time difference between the two vehicles entering the conflict zone [44]. Based on this, we define DGT. We generalize the conflict point to a conflict area—the intersection of the two bounding boxes’ swept areas, detected via the Separating Axis Theorem (SAT)—and calculate DGT as the time difference between the moments each vehicle first enters this area. A DGT of zero signifies a persistent overlap, prompting the removal of the shorter trajectory.
@@
In the temporal matching stage, we focused on time-series data. Each trajectory was synchronized with intersection entry/exit time and the corresponding traffic signal status. This temporal integration allowed for the fusion of supplementary information, such as traffic violations (e.g., red-light running), thus completing the comprehensive spatio-temporal profile for each TP.
@@
The full dataset is available through the figshare repository (in private currently). Figure 7 shows all the intersections, which are located in the central urban area. For privacy protection, the specific names and locations of these three intersections have been anonymized. Data was collected at these sites during clear daytime hours on selected days in January and May 2025. After an anomaly screening process, a total of approximately 5 hours of raw video footage was obtained. An overview of these three intersections is as follows:
@@
As the structure shown in Figure 8, the following components are available:
@@
  • Signal timings (signal): Manually annotated signal control data, sampled per second.
  • Flight log during videos (flightLog): Drone flight posture recording.
  • Maps (map): Georeferenced TIFF images and vector maps largely compatible with the Lanelet2 format [45].
  • Processed trajectories (traj/route/conflict): Offered in CSV format with a temporal resolution of 0.1 s (the annotations for conflicts and violations belong to different CSV files than the original tracks).
The full dataset is available through the figshare repository (in private currently). We wish to emphasize that FLUID proposes a processing framework rather than a definitive trajectory extraction result. The quality of the derived trajectories is intended as a baseline, open to further enhancement by the research community. We encourage researchers to apply state-of-the-art detection and tracking algorithms to the provided raw video footage and to explore various smoothing or filtering techniques for kinematics.
@@
Figure 9: Position and count comparison (Ground Truth/DFS/FLUID)
@@
Figure 9 presents the aligned trajectories and object count comparison. For consistency, all object classes were grouped into MV and VRU categories. The spatial comparison reveals that FLUID-extracted trajectories closely correspond with the DFS output. This positional accuracy is further corroborated by the RTS-GNSS ground truth: the Hausdorff distance between the reference and extracted trajectories for the test vehicle ranges $0\sim{0.97}m$, with errors under $0.3m$ on straight segments. Regarding object counts, DFS exhibited a $5\sim{6}\%$ miss rate against the ground-truth, a consequence of its policy to discard stationary and short trajectories. While this may improve tracking precision, it sacrifices recall. In contrast, FLUID achieved near-zero missed detections and a low ID switch rate of $2\sim{5}\%$, demonstrating performance comparable or superior to the DFS benchmark in comprehensive object detection.
@@
This observation is reinforced by examining individual speed profiles (the two MVs with the longest trajectories in the scene). As surveyed in Figure 11 (left), existing datasets exhibit inconsistent smoothing strategies. For instance, CitySim thresholds all near-zero speeds to zero, whereas SIND and inD apply this only to persistently stationary objects, leaving transient stops untreated. Smoothing granularities also vary widely, risking analytical biases. In contrast, FLUID enhances transparency by providing both raw and smoothed data (Figure 11, right). An analysis of the two longest vehicle trajectories again demonstrates that the 10 FPS processing yields accurate, stable speed profiles that align well with the DFS results, thus validating our methodological choices.
@@
To maintain a consistent comparative basis, our analysis is confined to MVs. We selected Time-to-Collision (TTC) as the primary predictive SSM, given its robustness after standardizing the temporal resolution of velocity data across datasets. Potential conflicts are initially identified using the minimum TTC (minTTC) observed between any two trajectories [44]. Subsequently, to improve precision, we introduce our Dynamic Gap Time (DGT) method for post-validation, which effectively filters out kinematically plausible but physically impossible conflicts (e.g., those separated by infrastructure). Drawing from a review of established practices [49], we define a conflict event using the thresholds: $DGT\leq{4.0s},TTC\leq{2.0s}$
@@
Figure 14: Per-cycle violation rates for straight and left-turn movements, calculated as a proportion of total MVs.
@@
Figure 15 shows the distribution of different entrance lanes and corresponding turns. It can be seen that, except for a very small number of vehicles with abnormal paths, the marking of turns is helpful in classifying different behavior patterns of traffic flow. Here, we provide some other related applications:
  • Driving Decision Modeling: Analyze the impact of interconnected information on driving decisions and traffic operations in interactive dilemmas involving two vehicles encountering conflicting directions at an intersection. [55]
  • Intersection Operation Quantification: Evaluate the safety and efficiency characteristics of different intersection control strategies. [26]
  • Conflict Correlation Analysis: For example, quantify conflict severity using PET or GT, and analyze the relationship between this severity and other kinematic parameters.[56]
  • Trajectory Generation: Learn from human mobility patterns to generate human-like and socially-inspired behaviors and movements. [57]
@@
Regarding raw data processing, the object detection and tracking code itself is based on the open-source YOLOv8 and SparseTrack. Therefore, the entire code is not repeated. The methods for post-processing the tracking results to obtain the FLUID dataset is publicly available on Github: https://github.com/sysu19351014/FLUID.

### v2

# captionstableheading

@@ Abstract
The trajectory data of traffic participants (TPs) is a fundamental resource for evaluating traffic conditions and optimizing policies, especially at urban intersections. Although data acquisition using drones is efficient, existing datasets still have limitations in scene representativeness, information richness, and data fidelity. This study introduces FLUID, comprising a fine-grained trajectory dataset that captures dense conflicts at typical urban signalized intersections, and a lightweight, full-pipeline framework for drone-based trajectory processing. FLUID covers three distinct intersection types, with approximately 5 hours of recording time and featuring over 20,000 TPs across 8 categories. Notably, the dataset records an average of 2.8 vehicle conflicts per minute across all scenes, with roughly 15 % of all recorded motor vehicles directly involved in these conflicts. FLUID provides comprehensive data, including trajectories, traffic signals, maps, and raw videos. Comparison with the DataFromSky platform and ground-truth measurements validates its high spatio-temporal accuracy. Through a detailed classification of motor vehicle conflicts and violations, FLUID reveals a diversity of interactive behaviors, demonstrating its value for human preference mining, traffic behavior modeling, and autonomous driving research.
@@
Methods for observing traffic participant (TP) mobility and interactions are primarily categorized into ego-centric, roadside, and aerial perspectives [7]. Ego-centric approaches, foundational to autonomous driving datasets like KITTI [8], Apolloscape [9], Waymo [10], and nuPlan [11], equip vehicles with multi-modal sensors to capture surrounding interactions [12], but are inherently limited by occlusions and a restricted field of view that prevent a complete scene understanding. Roadside methods, employing either single-sensor (e.g., NGSIM [13], Zen Traffic Data [14], I-24 MOTION [15], TJRD [16]) or multi-sensor systems (e.g., DLR-UT [17]), offer long-term observation but suffer from limited spatial coverage and potential error accumulation [18]. A critical drawback for both ground-based methods is that visible hardware can alter driver behavior, compromising data naturalness [19]. In contrast, aerial observation—typically using drones for their stability and cost-effectiveness—naturally overcomes these limitations by providing a complete, bird’s-eye view of the entire traffic scene and all simultaneous interactions within it. Crucially, high-altitude operation ensures the drone remains virtually unnoticed, thus preserving the naturalness of TP behavior. These advantages have spurred the development of numerous drone-based datasets for highways (e.g., highD [20], exiD [21], CQSkyEyeX [22], MiTra [23]), urban junctions (e.g., inD [24], SIND [25], Hohhot-HDI [26], Songdo Traffic [27], RounD [28]), and mixed urban environments (e.g., INTERACTION [29], pNEUMA [30], CitySim [31]).
Among the various traffic scenarios, urban intersections are particularly critical due to the significant efficiency losses and safety risks associated with the interrupted traffic flow. In the domain of intersection datasets, our review focuses on three key aspects: the captured scenes, the provided information, and the quality of trajectory data. Prominent existing datasets typically cover a limited number of scenarios (e.g., 3-4 intersections) [29, 24, 31, 25], but generally provide comprehensive information, including precise object dimensions and detailed scene elements. In contrast, the HDI dataset [26], while featuring diverse trajectory patterns across more intersections, lacks the aforementioned details. This discrepancy precludes a direct and fair comparison. Therefore, our subsequent analysis evaluates the datasets presented in Table 1 against these three aspects, with separate considerations for motor vehicles (MVs) and vulnerable road users (VRUs, including pedestrians and two-wheelers).
@@
Geometries and channelization Control methods MV arrival rate ($min^{-1}$) VRU arrival rate ($min^{-1}$) Conflict MV ratio Assoc. MV ($\leq$10 m) Structured Map MV types VRU types Conflict annot. Intention annot. Speed Position
INTERACTION 2019 FI, TI U (AWS) 34.29 1.82 1.80% 1.46 $\checkmark$ car pedestrian / bicycle
inD 2019 FI, TI U (RBL/prio.) 13.59 9.11 3.72% 1.25 $\checkmark$ car pedestrian, bicycle
pNEUMA 2020 Multiple Unknown 7.08 3.61 Car/Taxi, Medium/Heavy Vehicle.,Bus Motorcycle
CitySim 2022 FI, TI S (prot./perm.) / U (prio.) 27.69 8.14% 1.36 $\checkmark$ car
SIND 2022 FI S (perm.) 21.41 11.82 11.78% 1.74 $\checkmark$ car, truck, bus pedestrian, bicycle, motorcycle $\checkmark$
Songdo Traffic 2024 Multiple S (prot.) 2.03 0.06 3.12% 1.28 car/van, truck, bus motorcycle
FLUID (ours) 2025 FI, TI, FIDRT S (prot./perm.) 31.10 37.29 15.14% 1.66 $\checkmark$ car, tricycle, van, truck, bus, trailer pedestrian, moped $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
  • INTERACTION dataset/pNEUMA dataset/Songdo Traffic dataset: We use the publicly available multi-agent prediction version of INTERACTION dataset, which features pre-extracted interaction scenarios. For pNEUMA and Songdo Traffic, they do not have structured road network data, so we summarize the types of intersections by observing their sample images. Additionally, the statistical results of pNEUMA were calculated based on 60 major intersections.
  • FI/TI/FIDRT/Multiple: Abbreviations for intersection geometry and channelization types, as defined in the text. Multiple indicates that these types are included, and it also covers other abnormal intersections.
@@
  • Arrival rate: The average number of new MVs/VRUs arriving per minute during the recording period.
@@
The first aspect is the captured scenes, analyzed from both static (intersection type) and dynamic (traffic flow) perspectives.From a static perspective, intersections are classified by their geometries and channelization—such as three-way/four-way intersections (TI/FI), and four-way with dedicated right-turn lanes (FIDRT)—and by their control methods. The latter include unsignalized (e.g., uncontrolled/right-before-left, all-way-stop, priority-controlled) and signalized types, which feature permissive or protected phases for conflicting directions [32]. Different characteristics of intersections directly affect the speed and angle of TPs when crossing them, which is closely related to the situation of traffic conflicts. However, Figure 1 shows that existing drone datasets seldom focus on intersections with dedicated channelization and usually provide incomplete coverage of diverse control strategies. From a dynamic perspective, we evaluate traffic flow using three key metrics: TP arrival rates, MV conflict ratio, and the number of associated MV per conflict ($N_{CMVCP}$), where conflicts are identified using time-based surrogate safety measures (SSMs). A comparative analysis reveals several limitations in existing datasets, such as low or imbalanced participant counts(e.g., a low VRU proportion in INTERACTION and a low overall TP count in inD) and consistently low MV conflict ratio across several major datasets. Furthermore, the pNEUMA dataset’s high shooting altitude leads to decimeter-level target positioning accuracy, making it unsuitable for conflict analysis. In contrast, our proposed FLUID dataset shows clear advantages, featuring a higher overall TP arrival rate and a higher MV conflict ratio of over 15%. Its $N_{CMVCP}$ metric, comparable to that of SIND, further underscores its strength in capturing dense and interconnected traffic conflicts.
@@ Methods
The raw data was collected by a three-person team: one operator for the drone, and two observers who recorded traffic signal phases using ground-based cameras. We used a DJI Mini 3 drone to capture high-definition video at 4K resolution (3840×2160 pixels) and approximately 30 FPS (29.97 FPS). Due to regulations, the maximum altitude for drones is 120 meters, which is sufficient for capturing microscopic traffic behavior. The drone maintained a consistent flight altitude of $100\sim 120$ meters ($100\sim 105\;m$ for the FI scenario to capture finer VRU details; 120m120\;m for FIDRT and TI). The plane drift of this type of drone during stationary shooting does not exceed 1.5 meters, with a maximum recording duration per flight of up to 30 minutes. The precise flight altitude for each session can be retrieved from the drone’s flight logs. To synchronize the timestamps, the drone and ground cameras were aligned to a unified mobile phone clock with second-level precision, using the coordinated movement of a ground marker as a temporal reference.
@@
Due to the drone’s lightweight design, the raw footage was susceptible to wind-induced instability and required stabilization. We implemented a two-stage stabilization pipeline based on the open-source tools developed by Fonod et al. [35]. The lower level employed the feature detector based on video quality: AKAZE [36] is employed when sharpness remains consistent, leveraging its robust scale-space construction via non-linear diffusion filtering; BRISK [37] is preferred for videos with significant clarity variations. Both offer a favorable balance of speed and accuracy compared to alternatives like SIFT or ORB [38]. The upper level then utilized a RANSAC algorithm with block matching and motion compensation to robustly estimate inter-frame motion parameters while rejecting false matches from dynamic objects. This process effectively eliminated significant jitter from the footage. Moreover, we develop a masking program to define a Region of Interest (ROI) that strictly encompassed the intersection area. This step served to both anonymize areas outside the road network and reduce the computational load for subsequent processing. The FIDRT scenario was exempted from masking to preserve the full range of VRU activities. Finally, the videos were downsampled to 10 FPS. This frame rate is sufficient to capture the decision-making time range of human TPs and is also conducive to detecting the normal motion of VRUs [25]. The primary output of this stage is the set of pre-processed videos.
@@
Detection was accomplished using the YOLOv8 architecture [39], selected for its native support for both horizontal/oriented bounding boxes (HBB/OBB) and its efficiency as a single-stage detector. We prioritized an OBB representation as it provides a tighter and more accurate encapsulation of object boundaries and orientation from our aerial perspective, which is critical for the subsequent analysis of kinematic parameters and TP behaviors. To overcome single-dataset limitations and achieve higher detection accuracy across diverse object classes, we employ a multi-detector ensemble strategy. This strategy involves training the same YOLOv8 architecture independently on three distinct datasets, resulting in three sets of specialised model weights, each optimised to leverage the unique strengths of its training data.
@@
For each of the three training processes, the respective dataset was re-partitioned into training, validation, and test subsets using a 70%/20%/10% ratio, and each model instance was trained for 200 epochs. Subsequently, the final detections were generated through a category-based fusion of the outputs from these specialised detectors, as illustrated in Figure 4.
To evaluate the performance of each detector, we adopt the standard metrics from the YOLOv8 framework, including bounding box loss, classification loss, and Distribution Focal Loss (DFL) for both training and validation phases. The detection performance is quantified by Precision, Recall, and mean Average Precision (mAP) at Intersection over Union (IoU) thresholds of 0.5 and 0.5:0.95. Detailed definitions of each metric are provided in the supplementary materials [43][44]. F1-Confidence curve is a curve that shows the variation of F1-score as Confidence gradually increases. The F1-score exceeded 0.8 for all categories, indicating a well-balanced performance between precision and recall and confirming the robustness of the detection model.
In this process, We identified the object categories where each detector outperforms the others (termed its advantage categories). The final detection output for any given frame was constructed by integrating only the detections for these designated advantage categories from their respective specialist detector. This ensemble approach ensures a comprehensive and precise set of object detections.
Lightweight Tracking
@@
The most common time-based SSM, Post-Encroachment Time (PET), is ill-suited for this task because it requires a clear exit time from a conflict zone, which is ambiguous when considering full bounding boxes. We therefore turned to the concept of Gap Time (GT), which only considers the time difference between the two vehicles entering the conflict zone [49]. Based on this, we define DGT. We generalize the conflict point to a conflict area—the intersection of the two bounding boxes’ swept areas, detected via the Separating Axis Theorem (SAT)—and calculate DGT as the time difference between the moments each vehicle first enters this area. A DGT that remains zero for a sustained duration signifies a persistent overlap, prompting the removal of the shorter trajectory.
@@
In the temporal matching stage, we focused on time-series data. Each trajectory was synchronized with intersection entry/exit time and the corresponding traffic signal status. This temporal integration allowed for the fusion of supplementary information, such as traffic violations (e.g., red-light running).
@@
The full dataset is available through the figshare repository [50]. Figure 7 shows all the intersections, which are located in the central urban area. For privacy protection, the specific names and locations of these three intersections have been anonymized. Data was collected at these sites during clear daytime hours on selected days in January and May 2025. After an anomaly screening process, a total of approximately 5 hours of raw video footage was obtained. An overview of these three intersections is as follows:
@@
The formats of files in the FLUID dataset are: MPEG-4 Part 14 (MP4), Comma-Separated Values (CSV), Tagged Image File Format (TIFF), and OpenStreetMap (OSM). Since there are too many fields, their meanings are explained in the Markdown document README.md. As the structure shown in Figure 8, the following components are available:
@@
  • Signal timings (signal): Manually annotated signal control data in CSV format, sampled per second.
  • Flight log during videos (flightLog): Drone flight posture recording in CSV format.
  • Maps (map): Georeferenced TIFF images and vector maps largely compatible with the Lanelet2 (OSM) format [51].
  • Processed trajectories (traj/route/conflict): Offered in CSV format (the annotations for conflicts and violations belong to different files than the original tracks).
It is worth noting that FLUID offers not only a unified processing framework, but also high-quality trajectory extraction results, which act as a solid baseline and can be further enhanced by the research community.
@@
Figure 9: Position comparison (the coordinate of trajectories)
Table 2: Comparison of MV and VRU counts across different sources
Type Source Error
FLUID DFS Ground Truth (%)
MV 288 267 281 +2.5/-5.0
VRU 206 184 196 +5.1/-6.1
Total 494 451 477 +3.6/-5.4
@@
Figure 9 presents the aligned trajectories positions. For consistency, all object classes were grouped into MV and VRU categories. The spatial comparison reveals that FLUID-extracted trajectories closely correspond with the DFS output. This positional accuracy is further corroborated by the RTS-GNSS ground truth: the Hausdorff distance between the reference and extracted trajectories for the test vehicle ranges $0\sim{0.97}m$, with errors under $0.3m$ on straight segments. Table 2 reveals object counts. DFS exhibited a $5\sim{6}\%$ miss rate against the ground-truth, a consequence of its policy to discard stationary and short trajectories. While this may improve tracking precision, it sacrifices recall. In contrast, FLUID achieved near-zero missed detections and a low ID switch rate of $2\sim{5}\%$, demonstrating performance comparable or superior to the DFS benchmark in comprehensive object detection.
@@
This observation is reinforced by examining individual speed profiles (the two MVs with the longest trajectories in the FLUID scenario). In contrast, the left panel of Figure 11 illustrates that existing datasets employ inconsistent smoothing strategies. For instance, CitySim thresholds all near-zero speeds to zero, whereas SIND and inD apply this only to persistently stationary objects, leaving transient stops untreated. The variation in smoothing granularities across scenes—as seen in datasets like SIND and inD—is substantial and could lead to analytical biases. Conversely, FLUID enhances transparency by providing both raw and smoothed data (Figure 11, right). The analysis again demonstrates that the 10 FPS processing yields accurate, stable speed profiles that align well with the DFS results, thus validating our methodological choices.
@@
To maintain a consistent comparative basis, our analysis is confined to MVs. We selected Time-to-Collision (TTC) as the primary predictive SSM, given its robustness after standardizing the temporal resolution of velocity data across datasets. Potential conflicts are initially identified using the minimum TTC (minTTC) observed between any two trajectories [49]. Subsequently, to improve precision, we introduce our Dynamic Gap Time (DGT) method for post-validation, which effectively filters out kinematically plausible but physically impossible conflicts (e.g., those separated by infrastructure). Drawing from a review of established practices [55], we define a conflict event using the thresholds: ${0s}\leq{DGT}\leq{4.0s},{0s}\leq{TTC}\leq{2.0s}$.
@@
Figure 14: Per-cycle violation rates for straight and left-turn movements, calculated as a proportion of total MVs
@@
Benefiting from the fine-grained details of our dataset, we can categorize distinct traffic behavior patterns through turn labels, as illustrated in Figure 15. Beyond basic traffic flow analysis, the high precision of FLUID facilitates multi-domain research, including human preference mining, traffic behavior modeling, and autonomous driving. Figure 16 showcases three representative cases that highlight the dataset’s unique potential in these specialized research contexts.
@@
Figure 16: Unique application prospects of FLUID
Analysis of Passing and Yielding Behaviors.
Leveraging the standardized geometric layout, high density of passenger vehicles, and integrated traffic signal data in the FI scenario, FLUID provides a robust foundation for analyzing complex interaction behaviors. By calculating the convex hull of conflicting trajectories, we can accurately delineate conflict zones and analyze the temporal sequences of vehicles entering and exiting both these zones and the intersection boundaries. This methodology enables a refined quantification of passing and yielding dynamics for specific maneuvers. For instance, among the through-left interaction events identified during green phases across the FI dataset, we recorded 502 instances of left-turn yielding, 415 instances of through-moving yielding, and 5 instances where no clear yielding behavior was discernible (Figure 16, Left).
Spatiotemporal Violation Analysis of VRUs.
While previous research[61] (conducted on a scenario similar to FIDRT) was often constrained by limited tracking precision and necessitated grid-based spatial partitioning, the FLUID dataset provides individual-level trajectories of VRUs integrated with precise Lanelet2 semantic maps. Moreover, the significantly higher VRU arrival rate in FLUID, compared to existing benchmarks, allows for a more granular analysis of spatial occupancy. These features facilitate a deeper understanding of VRUs’ spatiotemporal violation intentions and movement patterns (Figure 16, Middle).
Optimization of Large Vehicle Detection
In the TI scenario, large vehicles constitute over 15% of the traffic, where their diverse dimensions present substantial challenges for accurate detection. By providing both raw data and bounding box labels, we offer significant optimization potential for large-scale target detection in complex environments. Ultimately, this provides a contribution of similar value to pNEUMA Vision [62] for multi-object tracking and detection research (Figure 16, Right).
In addition to these unique capabilities, our dataset also serves as a high-quality resource for broader applications in traffic engineering, such as:
  • Driving Decision Modeling: Analyze the impact of interconnected information on driving decisions and traffic operations in interactive dilemmas involving two vehicles encountering conflicting directions at an intersection [63].
  • Intersection Operation Quantification: Evaluate the safety and efficiency characteristics of different intersection control strategies [26].
  • Conflict Correlation Analysis: Quantify conflict severity using SSMs, and analyze the relationship between this severity and other kinematic parameters [64].
  • Trajectory Generation: Learn from human mobility patterns to generate human-like and socially-inspired behaviors and movements [65].
Data Availability
The full dataset is available at figshare [50]. The DOI to data record is 10.6084/m9.figshare.29974954.
@@
Regarding raw data processing, the object detection and tracking code itself is based on the open-source YOLOv8 and SparseTrack. Therefore, the entire code is not repeated. The methods for processing, verification and visualization of the FLUID dataset is publicly available on GitHub: https://github.com/sysu19351014/FLUID.